OpenStudy (anonymous):

how do youu solve for x in 2/3x + 1/2=5/6 ?

OpenStudy (smurfy14):

x=1/2

OpenStudy (anonymous):

2/3x = 2/6 6 = 3x x = 2

OpenStudy (smurfy14):

get x by itslef by subtracting 1/2 and then multiply each side by recipricol (3/2)

OpenStudy (anonymous):

Multiply all of the terms by 6. This will clear all of the fractions and leave 4x+3=5 Now solve for x x=1/2

OpenStudy (anonymous):

2/3x + 1/2=5/6 2/3x = 2/6 multiply both sides per 1/2 1/3x = 1/6 6 = 3x 2 = x so, the answer X = 2

OpenStudy (smurfy14):

corec youre wrong you multiply each side by 3/2 not 1/2

OpenStudy (anonymous):

Check your answer corec :)

OpenStudy (anonymous):

2/3x + 1/2=5/6 2/3X = 5/6 - 1/2 WE KNOW THAT 1/2 = 3/6 THEN 2/3X = 5/6 - 3/6 2/3X = 2/6 MULTIPLE BY 1/2 1/3X = 1/6 6 = 3X 2 =X TELL ME KNOW WHO IS WRONG? VICTORVECTOR

OpenStudy (anonymous):

how did youu get the 5 if youu multipled it by 6 and got 30 ? VICVEC

OpenStudy (smurfy14):

Brittbritt the answer is 1/2

OpenStudy (anonymous):

wait nevermined vicvec ii*figured it out LOl and ooh kayy smurfy ")

OpenStudy (anonymous):

thanks